OpenStudy (blow_pop2000):

karen read 24 pages in a book the book has a total of 63 pages what fraction of the book did karen read write the answer in simplest form a. 1/2 b. 2/5 c. 4/9 d. 9/4 help plz!!!!!!!

OpenStudy (maddy1251):

24/63 is what she read.

OpenStudy (maddy1251):

8/21. When I divided by 3.

OpenStudy (maddy1251):

It doesn't simplfy down any more than that.

OpenStudy (maddy1251):

Not sure which answer choice it would be then.

OpenStudy (maddy1251):

It cannot be D. That's impossible.

OpenStudy (maddy1251):

a. 1/2 b. 2/5 c. 4/9 d. 9/4 help plz!!!!!!! 8/21=0.38 So I guess we will find what fraction is closest to that. D is out. 4/9=0.44444444444... 2/5=0.4 1/2= 0.5 I could say B.

OpenStudy (superdavesuper):

i say typos; either in the Q itself or when it was posted ;)

OpenStudy (maddy1251):

Because 4/9 is showing a never ending 4. @superdavesuper I agree, or faulty answers.

OpenStudy (maddy1251):

The closest number, however, is 2/5.

OpenStudy (superdavesuper):

i will medal u twice 4 effort if i could @Maddy1251

OpenStudy (maddy1251):

Lol thanks :-) When I first looked at the problem I knew the answer wasn't up there.. So I am just assuming they want rounded or closest answer.

OpenStudy (blow_pop2000):

thx guys :) for ur help